There are two vehicles leaving for the same destination. One vehicle is 1.5 miles away and is traveling at 30mph. The other is 2.5 miles away and is traveling at 40mph. Which arrives first and by how much time?
D=RxT.

Found 3 solutions by Alan3354, mananth, sudhanshu_kmr:
Answer by Alan3354(69443)

1.5/30 = 0.05 hours
2.5/40 = 0.0625 hours
The closer is 1st by 0.0125 hours

Answer by mananth(16949)

1.5 miles away 30 mph
2.5 miles away 40 mph
..
first will take 1.5/30 *60 minutes = 3 minutes
second will take 2.5/40 *60 minutes = 3.75 minutes
..
you can conclude

Answer by sudhanshu_kmr(1152)

1.5 miles away , 30mph
time taken = (1.5/30 ) *60 minutes = 3 minutes
2.5 miles away, 40 mph
time taken = (2.5/40)*60 minutes = 3.75 minutes
so, first vehicle will come first.